Worked Out 65.3 grams of Lead(3) Sulfate Molar Conversions

Worked Out

65.3 grams of Lead(3) Sulfate

Molar Conversions
Answer #1

mass of Lead(3)sulfate = 65.3 grams

Lead(3)sulfate = Pb2(SO4)3 = 702.59 gram/mole

number of moles = mass of compound/molar mass

number of moles = 65.3/702.59 = 0.0929 molea

number of moles of Lead(3)sulfate present in 65.3 grams = 0.0929 moles.